Find the slope intercept form of the line passing through the point (-4,-9) and parallel to the line -2x-6y=7. I just dont need the answer i need how to do it as well.
Answer by Mathtut(3670)

you need to use the point slope formula which is
:
y-k=m(x-h) where m is the slope and (h,k) is any point on the line
:
we know that any line parallel to -2x-6y=7 has an equal slope. So the first step is to find the slope of this line
:
-2x-6y=7--->-6y=2x+7--->y=(-1/3)x-7/6....so the slope is -1/3
:
now we have the slope -1/3 and a point (-4,-9).....plug these numbers into our formula above
:
y-(-9)=(-1/3)(x-(-4))
:
y+9=(-1/3)(x+4)
:
y+9=-1/3x-4/3
:
y=(-1/3)x-31/3 slope intercept form